Find -: U(x1,x2)=16x1 + 8x2 axz O 16x1 + 8 O 16 + 8x2 O 24
Answer #1

Option 'C' is correct.

8 is the correct answer.

We are partially differentiating U with respect to x​​​​​​2​​​​​ so x​​1​​​ is constant.(hence partially differentiating 16x1 with respect to x​​​​​​2 will be 0.)

au/ar2 = a(16r1 + 8x2)/8x2 . = a(16r1)/3x2 + /8r2)/8x2 = 0+8 = 8 (Answer)

Therefore, Option 'C' is correct.